# references.bib

### v2

@@ Abstract
The Animal-AI Environment is a unique game-based research platform designed to facilitate collaboration between the artificial intelligence and comparative cognition research communities. In this paper, we present the latest version of the Animal-AI Environment, outlining several major new features that make the game more engaging for humans and more complex for AI systems. New features include interactive buttons, reward dispensers, and player notifications, as well as an overhaul of the environment’s graphics and processing for significant improvements in agent training time and quality of the human player experience. We provide detailed guidance on how to build computational and behavioural experiments with the Animal-AI Environment. We present results from a series of agents, including the state-of-the-art Deep Reinforcement Learning agent Dreamer-v3, on newly designed tests and the Animal-AI Testbed of 900 tasks inspired by research in the field of comparative cognition. The Animal-AI Environment offers a new approach for modelling cognition in humans and non-human animals, and for building biologically-inspired artificial intelligence.
@@ 1 Introduction
The Animal-AI Environment is the realisation of the idea that artificial intelligence (AI) research and cognitive science can and should work together to answer questions about cognition, behaviour, and intelligence. The Animal-AI Environment was first released in 2019 ([beyret2019animalai, crosby2019animal]), with a suite of (initially) held-out cognitive tasks (the Animal-AI Testbed) that were used to run The Animal-AI Olympics competition in 2020 ([crosby2020animal]). The environment is designed to offer a physically realistic virtual laboratory for building experiments inspired by research in comparative cognition, which studies and compares behaviour in non-human animals. In this paper, we present a major update to the environment, facilitating the development of a much wider range of behavioural experiments than was previously possible. The newest version of the Animal-AI Environment will interest researchers working on the development and evaluation of cognitively inspired AI, as well as comparative and developmental psychologists working on developing computational models of behaviour in humans and non-human animals.
The Animal-AI Environment has a simple structure, designed to emulate the classical set-up of psychological research on non-human animals. The agent is spawned in a large arena, emulating an enclosure or laboratory, and it can be controlled by either a computational model or a human player. The objective is to increase overall reward, by obtaining rewarding objects, while avoiding punishing ones within the time available. Rewarding objects correspond to appetitive stimuli such as food, while punishing objects correspond to aversive stimuli, such as electric shocks. The arena can be populated with a variety of other objects, such as ramps, blocks, tunnels, movable boxes, buttons, decoy goals, and dispensers. By using these objects as building blocks, an enormous variety of tasks can be generated, including setups similar to experiments from animal behaviour research ([crosby2020animal, voudouris2022evaluating]). Thousands of distinct tasks can be easily procedurally generated to build comprehensive test batteries ([crosby2020building]). The Animal-AI Environment is accompanied by several thousand open-source implementations of common experimental designs used in animal behaviour research, which can be freely used and adapted.
@@ 2 Collaborative Interdisciplinary Research with Animal-AI
Having a virtual laboratory for conducting behavioural tests on computational models encourages a virtuous cycle between AI and comparative cognition. Figure 1 presents a simplified schematic of this cycle. Researchers start by replicating a comparative cognition experiment in the Animal-AI Environment, ranging from tests of instrumental learning to tests of tool using and causal reasoning. They can then evaluate the performance of computational models on that test, either the baseline systems that we provide or those they design through the standardised Python interface. How these models perform on realistic, lab-like tasks informs researchers of their flaws or merits, encouraging the development of new hypotheses, theory, and experimental designs which can then be developed in the physical laboratory with non-human animals. These new designs can then be re-implemented in the Animal-AI Environment. Not only can this process improve and render more precise the theories and models used to explain animal cognition, but they can also inspire improvements in the architectures and methods of AI ([mitchener2022detect]).
@@
    A three-dimensional environment, with accurate, and fast physics.
@@
In contrast to comparative cognition and the cognitive sciences, there have been several new environments similar to the Animal-AI Environment developed in the AI community. Most similar to the Animal-AI Environment in being three-dimensional environments with realistic physics are MuJoCo ([todorov_2012]), DeepMind Lab ([beattie2016deepmind, leibo2018psychlab]), Minecraft-based environments (Malmö [johnson2016malmo], MineRL [guss2021minerl], MineDoJo [fan2022minedojo]), MiniWorld ([chevalier2024minigrid]), XLand ([humantimescaleadaptation2022]), Avalon ([albrecht2022avalon]), and ThreeDWorld ([gan2020threedworld, gan_open_2021]). The Animal-AI Environment distinguishes itself from these environments in being the only one targeted directly at the intersection between comparative cognition and AI, as well as being freely available, compatible with contemporary machine learning libraries in Python, facilitating customisable experimental designs, and being accompanied by an existing library of several thousand implementations of classic comparative cognition experiments (see Table 1). Similarly to XLand and ThreeDWorld, the Animal-AI Environment is built on top of Unity’s physics engine ([juliani2018unity]), providing fast real-world physics and three-dimensional graphics ([ward2020using]). We also make use of Unity’s open-source ml-agents Python API ([juliani2018unity]), which is integrated with existing reinforcement learning frameworks, such as Gymnasium ([towers_gymnasium_2023]), Dopamine ([castro2018dopamine]), and stable-baselines-3 ([stable-baselines3]). Unlike these environments and similarly rich environments such as Avalon, the Animal-AI Environment is more constrained in its scope. While these other environments are invaluable for the development of large, data-driven computational models, they often demand expensive hardware and large teams to use. The Animal-AI Environment is compact and simple enough to be used by small teams of AI researchers and cognitive scientists alike, while also scaling well to larger experiments. Nevertheless, Animal-AI’s physics and visuals are rich enough to facilitate complex experiments that are still too difficult for state-of-the-art systems ([crosby2020animal, voudouris2022direct, voudouris2024investigating]; see also Section 4.1.3).
@@
Most similar to the the Animal-AI Environment in scope and focus are DeepMind Lab, MiniWorld, and the Minecraft-based builds MineRL and Malmö. DeepMind Lab is an environment for training and testing autonomous agents on a range of tasks. It is designed to have convincing three-dimensional physics and simple visuals, similar to the Animal-AI Environment. It was designed to facilitate research on general and autonomous agents, and to build tests inspired by neuroscience. Indeed, DeepMind Lab has been used to study exploration and learning in young children, inspired by work in developmental psychology ([kosoy2020exploring]), as well as to improve reinforcement learning architectures ([leibo2018psychlab]). However, DeepMind Lab requires experiments to be designed in Lua, a language that is much less familiar to researchers than Python, which is how users interact with the Animal-AI Environment (see [chevalier2024minigrid]). Minecraft-based builds are similar to DeepMind Lab and the Animal-AI Environment in having a sophisticated physics engine and a definable ontology (i.e., blocks of different kinds that can populate the world). MiniWorld offers a three-dimensional environment with realistic physics for building customisable experiments for reinforcement learning systems. However, none of these environments are directly targeting comparative cognition research. The ontology of the Animal-AI Environment has been designed by comparative cognitive scientists to offer objects that are directly relevant to the methodologies of non-human animal behaviour research, including buttons, ramps, decaying and ripening rewards, and tools.
The Animal-AI Environment is accompanied by a library of experiments adapted from the comparative cognition literature, including 900 tasks testing spatial reasoning, working memory, numerosity, and tool-use ([crosby2020animal]), over 12,000 tasks testing object permanence ([voudouris2022evaluating]), and over 260,000 basic tasks for interacting with the environment. The environment has been used to develop more sophisticated reinforcement learning architectures ([lehuger2021fixed, mitchener2022detect, zakharov2021episodic]) as well as make progress on computational models of learning in non-human animals ([lind2021can, cardoso2023stimulus]). It also facilitates direct comparison to human participants, enabling the evaluation of the state-of-the-art in reinforcement learning ([burden2023inferring, voudouris2022direct, voudouris2024investigating]). The Animal-AI Environment is at the foundation of a rich and active research community at the intersection between comparative cognition and AI.
@@ 3 The Animal-AI Environment: A Comprehensive Overview
3.1 Environment Design and Specification
The Animal-AI Environment has two principle modes of use:
@@ 3.1 Environment Design and Specification
Substantial graphical improvements have been made to the Animal-AI Environment—not only cosmetic changes to improve the environment’s appearance (textured walls, animal skins, improved lighting) but also efforts to optimise environment rendering for faster performance. This includes pruning unnecessary casting and receiving of shadows; making use of the performance-orientated Universal Render Pipeline (URP); [^5] See https://unity.com/srp/universal-render-pipeline simplifying the shaders and lighting model used to render 3D objects; and reducing the resolution and anti-aliasing of textures and shadows. These optimisations complement efforts to create a more vibrant, engaging, and visually straightforward environmental appearance for human players, particularly children, while also increasing the frames per second (fps) at which the environment can run when using camera-based observations (see Section 3.5).
3.2 The Arena and Objects
@@ 3.2.1 The Arena & Agent
As in earlier versions of the environment, the arena is a 40$\times$40 unit space (where the agent is a sphere with a diameter of 1), including a tiled floor and four surrounding walls resembling white picket fences. [^6] Note that there is no explicit height restriction on the arena. The Arena object can be configured with three parameters. The reward for an episode begins at 0 and decreases in increments of $\frac{1}{t}$ per timestep, increasing when the agent interacts with a positively valenced object (see 3.2.4) and decreasing upon interaction with a negatively valenced object. The arena pass mark, passMark, allows the researcher to define the reward threshold that should constitute a ‘pass’ on that particular episode, providing an easy metric for success or failure across episodes. It also determines the notifications (if turned on) that players receive at the end of an episode (see Section 3.6). The arena time, timeLimit, provides the total number of time steps an agent or user can take before the episode ends. If timeLimit is set to [0], then the reward value no longer decreases with each time step.
The agent is of size 1$\times$1$\times$1 and can spawn anywhere in the Arena. It can take one of three forms: a hedgehog, a panda, or a pig, enabling the creation of an engaging narrative for studies with children (see Figure 3). The agent can be frozen at the start of an episode for a specified number of steps, using the frozenAgentDelays parameter, a new feature to the Animal-AI Environment. There is no reward decrement during the frozen period and all other objects continue to move and interact. This allows the agent to observe occurrences in the environment before being allowed to move.
@@ 3.2.2 Immovable Objects
  • Transparent Tunnels: A non-opaque hemi-cylindrical tunnel used as an obstacle to be navigated through and/or around.
@@ 3.2.3 Movable Objects
There are six movable objects in the Animal-AI Environment: hollow-box, light and heavy rectangular blocks, U-shaped, L-shaped, and J-shaped blocks, all of which can be pushed around. These blocks can be used as tools for moving other objects (other movable objects or any of the valenced rewards discussed below). These objects can be resized, rotated, and re-positioned freely, but their colour cannot be changed. In the Animal-AI Environment, unlike in earlier versions, blocks do not have a “cardboard box”- or “wood”- style texture, instead being uniform grey in colour with light grey trims along the vertices.
@@ 3.5 Using Animal-AI with Computational Models
The Animal-AI Environment has a straightforward Python API for integration with computational models, based on the ml-agents library ([juliani2018unity]). The environment returns information about the state of the environment at each time step, including the agent’s first-person observations, position, velocity, and current reward, and the model must return one action from the set of permissible actions in order to evolve the environment. Thus, any model implemented in Python that returns single actions in response to observations can be integrated into the Animal-AI Environment.
@@ 3.7 Designing Experiments in Animal-AI
We also provide a configuration design assistant, which allows users to visualise, edit, and create configuration files through a graphical web application written in Python. [^9] https://github.com/Kinds-of-Intelligence-CFI/aai-config-assist The assistant facilitates task building by removing the need to manually write YAML files and by giving users immediate visual feedback about the configuration being designed. The assistant provides considerable interactivity with many application callbacks. These include moving, resizing, deleting, and adding items to an Animal-AI arena as well as fully designing a new arena from scratch.
@@
Figure 10 shows the configuration design assistant. Both represent the same YAML configuration. Users can place new objects in the environment with definable positions, rotations, sizes, and colours, and the assistant detects any overlapping objects. When a configuration contains such an overlap, the problematic items are spawned in separate locations according to the precedence set out in the configuration file. This is exemplified in the right panel, where a ramp is spawned in the bottom left portion of the arena rather than to the right of the central blue wall as displayed in the configuration assistant representation. The assistant signals the potential overlap with a red border around the object. New items can effortlessly be added and supported by the assistant as the environment develops. The assistant is lightweight and quicker to launch than Animal-AI, thus streamlining configuration design. It has been used by both in-house and external researchers to speed up the experimental design process.
@@
Figure 10: The Animal-AI configuration design companion tool. Left: The graphical user interface for the tool, built in Python. Right: The rendered configuration in the Animal-AI Environment.
@@
Task configurations can also be procedurally generated. A simple accompanying command line tool is provided to design and generate precisely specified variations of a task. [^10] https://github.com/wschella/procgen-companion Each generated file is a valid Animal-AI configuration. The tool supports selecting random colours, specifying a fixed list of options, conditional values dependent on other choices already made, labels, and much more. Procedural generation allows for the quick creation of large test batteries, making behavioural evaluation more robust. For instance, researchers can study whether the performance of agents is sensitive to changes that ought to be irrelevant to success on the task, such as the colours of some objects.
@@
To demonstrate the ease with which one can design and run experiments in the Animal-AI Environment, we present three experiments. All coordinates are presented as ($x$,$y$,$z$) where $y$ is the vertical axis. The first experiment examines goal-directed navigation using a simple foraging task. The agent is spawned at a random $x$-coordinate at ($x$,0,5), with a random initial rotation. At the centre of the arena (20,0,20) is a SpawnerTree, which spawns 10 GoodGoalMulti goals of size 2. The agent must obtain them all before their health runs out.
@@
The second experiment is based on the well-known operant chamber experiments ([heron1939apparatus]) that have been used to study learning in a wide variety of animals ([macphail1982vertebrates]). In the classic design, the animal is placed in a chamber with a lever which can be pressed to return a food reward. Naive animals from which food has been withheld for a few hours are placed in the chamber, and learn through trial-and-error to operate the lever in return for rewards. An important feature of this task design in early work on animal behaviour was that lever presses were automatically recorded. Since the operant chamber is straightforward to design, several animals could be studied in parallel. We recreate a version of an operant chamber in the Animal-AI Environment by using the SpawnerButton object placed at (38.5,0,18.8), which, when pressed, spawns a GoodGoal next to it at (39,3,22). The button is enclosed by walls on three sides to ensure that the agent approaches it from the front. The agent is spawned at a random position in the arena. To succeed, the agent must press the button and then navigate to obtain the reward. We also construct a training curriculum, in line with work in reinforcement learning ([narvekar2020curriculum]) and on shaping in animal behaviour theory ([skinner1958reinforcement, peterson2004day]). The curriculum introduces the agent to the button and the consequences of pressing it. First, the agent is spawned at (36,0,20), directly in front of the bottom, such that it is likely to hit it by accident. Then, it is spawned further away, at (30,0,20), and then at (20,0,20). In the fourth stage, it is spawned at (20,0,$z$), where $z$ is randomly selected. Finally, it is presented with the final task, in which it is spawned at a random ($x$,$z$) coordinate. By comparing performance with and without the curriculum, its effect on learning can be directly assessed.

### v3

@@ Abstract
The Animal-AI Environment is a unique game-based research platform designed to facilitate collaboration between the artificial intelligence and comparative cognition research communities. In this paper, we present the latest version of the Animal-AI Environment, outlining several major features that make the game more engaging for humans and more complex for AI systems. These features include interactive buttons, reward dispensers, and player notifications, as well as an overhaul of the environment’s graphics and processing for significant improvements in agent training time and quality of the human player experience. We provide detailed guidance on how to build computational and behavioural experiments with the Animal-AI Environment. We present results from a series of agents, including the state-of-the-art deep reinforcement learning agent Dreamer-v3, on newly designed tests and the Animal-AI Testbed of 900 tasks inspired by research in the field of comparative cognition. The Animal-AI Environment offers a new approach for modelling cognition in humans and non-human animals, and for building biologically inspired artificial intelligence.
@@ 1 Introduction
The Animal-AI Environment is the realisation of the idea that artificial intelligence (AI) research and cognitive science can and should work together to answer questions about cognition, behaviour, and intelligence. The Animal-AI Environment was first released in 2019 ([beyret2019animalai, crosby2019animal]), with a suite of (initially) held-out cognitive tasks (The Animal-AI Testbed) that were the basis of the The Animal-AI Olympics competition in 2020 ([crosby2020animal]). The environment is designed to offer a physically realistic virtual laboratory for building experiments inspired by research in comparative cognition, the field that studies and compares behaviour in non-human animals. In this paper, we present a major update to the environment, facilitating the development of a much wider range of behavioural experiments than was previously possible. The newest version of the Animal-AI Environment will interest researchers working on the development and evaluation of cognitively inspired AI, as well as comparative and developmental psychologists working on developing computational models of behaviour in humans and non-human animals.
The Animal-AI Environment has a simple structure, designed to emulate the classical set-up of psychological research on non-human animals. The agent is spawned in a large arena, emulating an enclosure or laboratory, and it can be controlled by either a computational model or a human player. The objective is, within a specified time limit, to increase overall reward by obtaining rewarding objects and avoiding punishing ones. Rewarding objects correspond to appetitive stimuli such as food, while punishing objects correspond to aversive stimuli, such as electric shocks. The arena can be populated with a variety of other objects, such as ramps, blocks, tunnels, movable boxes, buttons, decoy goals, and dispensers. By using these objects as building blocks, an enormous variety of tasks can be generated, including setups similar to experiments from animal behaviour research ([crosby2020animal, voudouris2022evaluating]). Thousands of distinct tasks can be easily procedurally generated to build comprehensive test batteries ([crosby2020building]). The Animal-AI Environment is accompanied by several thousand open-source implementations of common experimental designs used in animal behaviour research, which can be freely used and adapted.
@@ 2 Collaborative Interdisciplinary Research with Animal-AI
Having a virtual laboratory for conducting behavioural tests on computational models encourages a virtuous cycle between AI and comparative cognition. Figure 1 presents a simplified schematic of this cycle. Researchers start by replicating a comparative cognition experiment in the Animal-AI Environment, ranging from tests of instrumental learning to tests of tool use and causal reasoning. They can then evaluate the performance of computational models on that test, either the baseline systems that we provide or those they design through the standardised Python interface. How these models perform on realistic, lab-like tasks informs researchers of their flaws or merits, encouraging the development of new hypotheses, theories, and experimental designs which can then be developed in the physical laboratory with non-human animals. These new designs can then be re-implemented in the Animal-AI Environment. Not only can this process improve and render more precise the theories and models used to explain animal cognition, but it can also inspire improvements in the architectures and methods of AI ([freire2024sequential, mitchener2022detect]).
@@
    A three-dimensional environment, with accurate and fast physics.
@@
In contrast to comparative cognition and the cognitive sciences, there are several environments similar to the Animal-AI Environment that have been developed in the AI community. Most similar to the Animal-AI Environment in being three-dimensional environments with realistic physics are MuJoCo ([todorov_2012]), DeepMind Lab ([beattie2016deepmind, leibo2018psychlab]), Minecraft-based environments (Malmö [johnson2016malmo], MineRL [guss2021minerl], MineDoJo [fan2022minedojo]), MiniWorld ([chevalier2024minigrid]), XLand ([humantimescaleadaptation2022]), Avalon ([albrecht2022avalon]), and ThreeDWorld ([gan2020threedworld, gan_open_2021]). The Animal-AI Environment distinguishes itself from these environments in being the only one targeted directly at the intersection between comparative cognition and AI, as well as being freely available, compatible with contemporary machine learning libraries in Python, facilitating customisable experimental designs, and being accompanied by an existing library of several thousand implementations of classic comparative cognition experiments (see Table 1). Similarly to XLand and ThreeDWorld, the Animal-AI Environment is built on top of Unity’s physics engine ([juliani2018unity]), providing fast real-world physics and three-dimensional graphics ([ward2020using]). We also make use of Unity’s open-source ml-agents Python API ([juliani2018unity]), which is integrated with existing reinforcement learning frameworks, such as Gymnasium ([towers_gymnasium_2023]), Dopamine ([castro2018dopamine]), and stable-baselines-3 ([stable-baselines3]). Unlike these environments and similarly rich environments such as Avalon, the Animal-AI Environment is more constrained in its scope. While these other environments are invaluable for the development of large, data-driven computational models, they often demand expensive hardware and large research teams. The Animal-AI Environment is compact and simple enough to be used by small teams of AI researchers and cognitive scientists alike, while also scaling well to larger experiments. Nevertheless, Animal-AI’s physics and visuals are rich enough to facilitate complex experiments that are still too difficult for state-of-the-art systems ([crosby2020animal, voudouris2022direct, voudouris2024investigating]; see also Section 4.1.3).
@@
Most similar to the Animal-AI Environment in scope and focus are DeepMind Lab, MiniWorld, and the Minecraft-based builds MineRL and Malmö. DeepMind Lab is an environment for training and testing autonomous agents on a range of tasks. It is designed to have convincing three-dimensional physics and simple visuals, similar to the Animal-AI Environment. It was designed to facilitate research on general and autonomous agents, and to build tests inspired by neuroscience. Indeed, DeepMind Lab has been used to study exploration and learning in young children, inspired by work in developmental psychology ([kosoy2020exploring]), as well as to improve reinforcement learning architectures ([leibo2018psychlab]). However, DeepMind Lab requires experiments to be designed in Lua, a language that is much less familiar to researchers than Python, which is how users interact with the Animal-AI Environment (see [chevalier2024minigrid]). Minecraft-based builds are similar to DeepMind Lab and the Animal-AI Environment in having a sophisticated physics engine and a definable ontology (i.e., blocks of different kinds that can populate the world). MiniWorld offers a three-dimensional environment with realistic physics for building customisable experiments for reinforcement learning systems. However, none of these environments are directly targeting comparative cognition research. The ontology of the Animal-AI Environment has been designed by comparative cognitive scientists to offer objects that are directly relevant to the methodologies of non-human animal behaviour research, including buttons, ramps, decaying and ripening rewards, and tools.
The Animal-AI Environment is accompanied by a library of experiments adapted from the comparative cognition literature, including 900 tasks testing spatial reasoning, working memory, numerosity, and tool-use ([crosby2020animal]), over 12,000 tasks testing object permanence ([voudouris2022evaluating]), over 5,000 tasks testing object affordances ([rutar4924246general]), and over 260,000 basic tasks for interacting with the environment. The environment has been used to develop more sophisticated reinforcement learning architectures ([freire2024sequential, lehuger2021fixed, mitchener2022detect, zakharov2021episodic]) as well as make progress on computational models of learning in non-human animals ([lind2021can, cardoso2023stimulus]). It also facilitates direct comparison to human participants, enabling the evaluation of the state-of-the-art in reinforcement learning ([burden2023inferring, voudouris2022direct, voudouris2024investigating]) and multi-modal large language models ([mecattafslater2024lessconversation]). The Animal-AI Environment is therefore at the foundation of a rich and active research community at the intersection of comparative cognition and AI.
@@ 3 The Animal-AI Environment: A Comprehensive Overview
3.1 Environment Design and Modes of Use
The Animal-AI Environment has two principal modes of use:
@@ 3.1 Environment Design and Modes of Use
Substantial graphical improvements have been made to the Animal-AI Environment—not only cosmetic changes to improve the environment’s appearance (textured walls, animal skins, improved lighting) but also efforts to optimise environment rendering for faster performance. This includes pruning unnecessary casting and receiving of shadows; making use of the performance-orientated Universal Render Pipeline (URP); [^3] See https://unity.com/srp/universal-render-pipeline. simplifying the shaders and lighting model used to render 3D objects; and reducing the resolution and anti-aliasing of textures and shadows. These optimisations complement efforts to create a more vibrant, engaging, and visually straightforward environmental appearance for human players, particularly children, while also increasing the frames per second (fps) at which the environment can run when using camera-based observations (see Section 3.5).
3.2 Environment Specification
@@ 3.2.1 The Arena & Agent
As in earlier versions of the environment, the arena is a 40$\times$40 unit space (where the agent is a sphere with a diameter of 1), including a tiled floor and four surrounding walls resembling white picket fences. [^4] Note that there is no explicit height restriction on the arena. The Arena object can be configured with three parameters. The arena time, timeLimit, provides the total number of time steps an agent or user can take before the episode ends. The reward for an episode begins at 0 and decreases in increments of $\frac{1}{t}$ per timestep, where t is the value of timeLimit, increasing when the agent interacts with a positively valenced object (see 3.2.4) and decreasing upon interaction with a negatively valenced object. If timeLimit is set to 0, then the reward value no longer decreases with each time step. The arena pass mark, passMark, allows the researcher to define the reward threshold that should constitute a ‘pass’ on that particular episode, providing an easy metric for success or failure across episodes. It also determines the notifications (if turned on) that players receive at the end of an episode (see Section 3.6).
The agent is of size 1$\times$1$\times$1 and can spawn anywhere in the Arena. It can take one of three forms: a hedgehog, a panda, or a pig, enabling the creation of an engaging narrative for studies with children (see Figure 3). The agent can be frozen at the start of an episode for a specified number of steps, using the frozenAgentDelays parameter, a new feature in the Animal-AI Environment. There is no reward decrement during the frozen period and all other objects continue to move and interact. This allows the agent to observe occurrences in the environment before being allowed to move.
@@ 3.2.2 Immovable Objects
  • Transparent Tunnels: A non-opaque hemicylindrical tunnel used as an obstacle to be navigated through and/or around.
@@ 3.2.3 Movable Objects
There are six movable objects in the Animal-AI Environment: hollow boxes, light and heavy rectangular blocks, U-shaped, L-shaped, and J-shaped blocks, all of which can be pushed around. These blocks can be used as tools for moving other objects (other movable objects or any of the valenced rewards discussed below). These objects can be resized, rotated, and re-positioned freely, but their colour cannot be changed. In the Animal-AI Environment, unlike in earlier versions, blocks do not have a “cardboard box”- or “wood”- style texture, instead being uniform grey in colour with light grey trims along the vertices.
@@ 3.5 Using Animal-AI with Computational Models
The Animal-AI Environment has a straightforward Python API for integration with computational models, based on the ml-agents library ([juliani2018unity]). The environment returns information about its state at each time step, including the agent’s first-person observations, position, velocity, and current reward, and the model must return one action from the set of permissible actions in order to evolve the environment. Thus, any model implemented in Python that returns single actions in response to observations can be integrated into the Animal-AI Environment.
@@ 3.7 Designing Experiments in Animal-AI
We also provide a configuration design assistant, which allows users to visualise, edit, and create configuration files through a graphical web application written in Python. [^7] See https://github.com/Kinds-of-Intelligence-CFI/aai-config-assist. The assistant facilitates task building by removing the need to manually write YAML files and by giving users immediate visual feedback about the configuration being designed. The assistant provides considerable interactivity with many application callbacks. These include moving, resizing, deleting, and adding items to an Animal-AI arena as well as fully designing a new arena from scratch.
@@
Figure 10 shows the configuration design assistant. Both represent the same YAML configuration. Users can place new objects in the environment with definable positions, rotations, sizes, and colours, and the assistant detects any overlapping objects. When a configuration contains such an overlap, the problematic items are spawned in separate locations according to the precedence set out in the configuration file. This is exemplified in the right panel, where a ramp is spawned in the bottom left portion of the arena, rather than to the right of the central blue wall as displayed in the configuration assistant representation. The assistant signals the potential overlap with a red border around the object. New items can effortlessly be added and supported by the assistant as the environment develops. The assistant is lightweight and quicker to launch than Animal-AI, thus streamlining configuration design. It has been used by by researchers to speed up the experimental design process.
@@
Figure 10: The Animal-AI configuration design companion tool. Left: The graphical user interface for the tool, built in Python. Right: The rendered configuration in the Animal-AI Environment. Note the displaced grey ramp in the bottom left of the arena, as a result of a detected clash in the configuration tool.
@@
Task configurations can also be procedurally generated. A simple accompanying command line tool is provided to design and generate precisely specified variations of a task. [^8] See https://github.com/wschella/procgen-companion. Each generated file is a valid Animal-AI configuration. The tool supports selecting random colours, specifying a fixed list of options, conditional values dependent on other choices already made, labels, and much more. Procedural generation allows for the quick creation of large test batteries, making behavioural evaluation more robust. For instance, researchers can study whether the performance of agents is sensitive to changes that ought to be irrelevant to success on the task, such as the colours of some objects.
@@
To demonstrate the ease with which one can design and run experiments in the Animal-AI Environment, we present three experiments. All coordinates are presented as ($x$,$y$,$z$) where $y$ is the vertical axis. The first experiment examines goal-directed navigation using a simple foraging task, emulating experiments of foraging behaviour in a wide variety of animals (see, e.g., [pyke1977optimal], [schoener1971theory]). The agent is spawned at a random $x$-coordinate at ($x$,0,5), with a random initial rotation. At the centre of the arena (20,0,20) is a SpawnerTree, which spawns 10 GoodGoalMulti goals of size 2. The agent must obtain them all before their health runs out. The agent maximises reward by finding the shortest path between the spawned goals, known as foraging by energy maximisation ([king2022optimal, pyke1977optimal]).
@@
The second experiment is based on the well-known operant chamber experiments ([heron1939apparatus]) that have been used to study learning in a wide variety of animals ([macphail1982vertebrates]). In the classic design, the animal is placed in a chamber with a lever which can be pressed to return a food reward. Naive animals from which food has been withheld for a few hours are placed in the chamber, and learn through trial-and-error to operate the lever in return for rewards. An important feature of this task design in early work on animal behaviour was that lever presses were automatically recorded. Since the operant chamber is straightforward to design, several animals could be studied in parallel, a property mirrored by the Animal-AI Environment for computational models. We recreate a version of an operant chamber in the Animal-AI Environment by using the SpawnerButton object placed at (38.5,0,18.8), which, when pressed, spawns a GoodGoal next to it at (39,3,22). The button is enclosed by walls on three sides to ensure that the agent approaches it from the front. The agent is spawned at a random position in the arena. To succeed, the agent must press the button and then navigate to obtain the reward. We also construct a training curriculum, in line with work in reinforcement learning ([narvekar2020curriculum]) and on shaping in animal behaviour theory ([skinner1958reinforcement, peterson2004day]). The curriculum introduces the agent to the button and the consequences of pressing it. First, the agent is spawned at (36,0,20), directly in front of the bottom, such that it is likely to hit it by accident. Then, it is spawned further away, at (30,0,20), and then at (20,0,20). In the fourth stage, it is spawned at (20,0,$z$), where $z$ is randomly selected. Finally, it is presented with the final task, in which it is spawned at a random ($x$,$z$) coordinate. By comparing performance with and without the curriculum, its effect on learning can be directly assessed.